### 30 Dec 2023 v1

@@ 7.3 Results in the Ant Maze
To investigate how effective POSG is across various environments, we evaluated our proposed POSG algorithm on the hard-exploration Ant Maze task. It turns out that our approach significantly outperforms other baseline methods in terms of control performance. As shown in Fig. 6, only our method can reach the goal and learn the optimal policy in the hard-exploration ant maze. Other baseline methods cannot find the sparse reward even if the demonstrations are provided. This result demonstrates that our POSG can better exploit sparse-reward demonstrations by merging the demonstrations’ distribution information and reward signals compared with other RLfD methods. Fig. 6 shows the changing curves of the MMD distance in the Ant Maze task. Our POSG method can efficiently decrease the distance between the training policy and sparse-reward demonstrations. Hence, POSG learns expert-like behaviors with demonstrations quickly. In contrast, the PPO+D agent cannot effectively learn the behaviors of demonstrations. Due to the sparsity of the reward function, the agent’s policy often deteriorates during the policy optimization process, and the agent always falls near the initial point at the end of training.
Figure 6: (a) Learning curves of success rate on the Ant Maze task; (b) MMD distance changing trends of POSG and PPO+D.
@@ 7.4.1 Sensitivity to Demonstrations
Figure 7: (a) Learning curves of success rate on the Ant Maze task; (b) MMD distance changing trends of POSG and PPO+D.
Demonstrations with different amounts. We selected five demonstration datasets with different amounts for comparison on the SparseHopper task. Notice that the POSG experiment is conducted with six trajectories of 500 state-action pairs as demonstrations. The black dashed line in Fig. 7 represents the expert-level performance. In Fig. 7, The $y$-value corresponding to the dashed line is the value of the average reward per step learned by the PPO algorithm under the ground-truth reward setting. Fig. 7 presents the result of this ablation experiment. Our POSG algorithm produces nearly identical results with different numbers of sparse-reward demonstrations. This result suggests that our POSG algorithm can achieve impressive performance using only a few demonstrations, even a single trajectory. Hence, our method decreases the requirement of demonstration quantity, which facilitates the application of the algorithm in real-world scenarios.
Demonstrations with different qualities. We obtained the demonstration datasets with different qualities by generating and storing the trajectory data in the different training phases. The corresponding result is shown in Fig. 8. In Fig. 8, the black dashed line represents the expert-level performance, and the gray dashed line represents the medium-level performance. The dashed line in Fig. 8 has the same meaning as that in Fig. 7. The average return of POSG increases as the proportion of high-quality data increases. This result indicates that the quality of demonstrations will significantly affect the performance of POSG and that high-quality demonstration data can contribute to policy optimization to some extent.
Figure 8: (a) Learning curves of success rate on the Ant Maze task; (b) MMD distance changing trends of POSG and PPO+D.
@@ 8 Conclusion
This article proposes a novel RLfD algorithm called POSG that realizes efficient exploration and policy optimization by introducing a smooth guidance reward. More specifically, we obtain the importance of a trajectory by considering the distributional distance between policies and the returns of the associated trajectories. Then, the smooth guidance reward is computed by smoothly averaging the trajectory importance over the trajectory space. We theoretically prove that smooth guidance rewards cause a higher performance improvement bound and deduce a new worst-case lower bound on performance improvement. We conducted various experiments by benchmarking POSG with several state-of-the-art RL algorithms in four environments. Experimental results demonstrate POSG’s superiority over other baseline approaches in terms of performance and convergence speed.
Declarations
  • Funding This work was supported by the National Natural Science Foundation of China Grant (Nos. 62141605)
  • Conflict of interest The authors declare that there is no conflict of interest.
  • Availability of data and materials The datasets generated during and/or analyzed during the current study are available from the corresponding author upon reasonable request.
  • Authors’ contributions Guojian Wang: Conceptualization, Methodology, Software, Formal analysis, Writing–Original Draft. Faguo Wu: Writing–Review and Editing, Supervision. Xiao Zhang: Validation, Supervision, Funding acquisition. Tianyuan Chen: Formal analysis, Software, Visualization.
@@ Appendix A POSG Algorithm Training Process in Discrete Spaces
1:learning rate $\alpha$, offline sparse-reward demonstration dataset $\mathcal{P}$, policy update frequency $K$
2:Initialize policy weights $\theta$ of each agent.
3:// Buffer that stores key-value pairs of a state-action tuple $(\boldsymbol{s},\boldsymbol{a})$ and a list of values of trajectory importance that include $(\boldsymbol{s},\boldsymbol{a})$
4:Initialize $\mathcal{M}\leftarrow\emptyset\ \forall(\boldsymbol{s},\boldsymbol{a})$
5:for $\text{episode}\in\{0,\dots,T\}$ do
6:     $\tau\leftarrow\emptyset$    // Store state-action pairs for current episode
7:     $R\leftarrow 0$    // Accumulate rewards for current episode
8:     for each step in $\{1,\dots,T\}$ do
9:         Choose $\boldsymbol{a}$ from $\boldsymbol{s}$ using $\pi_{\theta}$
10:         take action $\boldsymbol{a}$ and observe $\boldsymbol{r}_{e}$ and $\boldsymbol{s}^{\prime}$
11:         $\tau\leftarrow\tau\cup\{(\boldsymbol{s},\boldsymbol{a})\}$; $R\leftarrow R+\boldsymbol{r}_{e}$
12:     end for
13:     Compute the trajectory weight $\omega(\tau|\mathcal{M}_{E})$ and the joint return $R_{j}(\tau)$
14:     Compute the trajectory importance $I(\tau)=\omega(\tau|\mathcal{M}_{E})R_{j}(\tau)$
15:     for each $((\boldsymbol{s},\boldsymbol{a}))$ in $\tau_{e}$ do
16:         $\boldsymbol{r}_{i}=\mathbb{E}_{I\sim\mathcal{M}(\boldsymbol{s},\boldsymbol{a})}[I]$  // Compute guidance rewards for each state-action pair
17:     end for
18:     if $\text{episode}\,\%\,K==0$ then
19:         Compute $A_{e}$ using environmental rewards
20:         Compute $A_{i}$ using guidance rewards
21:         $\theta\leftarrow\theta+\alpha\nabla\eta_{e}(\theta)$   // Update parameters using environmental rewards
22:         $\theta\leftarrow\theta+\alpha\nabla\eta_{i}(\theta)$     // Update parameters using guidance rewards
23:     end if
24:end for
@@ Appendix B POSG Algorithm Training Process in High-Dimensional Continuous Spaces
1:learning rate $\alpha$, offline sparse-reward demonstration dataset $\mathcal{P}$, policy update frequency $K$
2:Initialize policy weights $\theta$ of each agent.
3:// Buffer that stores key-value pairs of a state-action tuple $(\boldsymbol{s},\boldsymbol{a})$ and a list of values of trajectory importance that include $(\boldsymbol{s},\boldsymbol{a})$
4:Initialize $\mathcal{M}\leftarrow\emptyset\ \forall(\boldsymbol{s},\boldsymbol{a})$
5:for $\text{episode}\in\{0,\dots,T\}$ do
6:     $\tau\leftarrow\emptyset$    // Store state-action pairs for current episode
7:     $R\leftarrow 0$    // Accumulate rewards for current episode
8:     for each step in $\{1,\dots,T\}$ do
9:         Choose $\boldsymbol{a}$ from $\boldsymbol{s}$ using $\pi_{\theta}$
10:         take action $\boldsymbol{a}$ and observe $\boldsymbol{r}_{e}$ and $\boldsymbol{s}^{\prime}$
11:         $\tau\leftarrow\tau\cup\{(\boldsymbol{s},\boldsymbol{a})\}$; $R\leftarrow R+\boldsymbol{r}_{e}$
12:     end for
13:     Compute the trajectory weight $\omega(\tau|\mathcal{M}_{E})$ and the joint return $R_{j}(\tau)$
14:     Compute the guidance reward $I(\tau)=\omega(\tau|\mathcal{M}_{E})R_{j}(\tau)$
15:     if $\text{episode}\,\%\,K==0$ then
16:         Compute $A_{e}$ using environmental rewards
17:         Compute $A_{i}$ using guidance rewards
18:         $\theta\leftarrow\theta+\alpha\nabla\eta_{e}(\theta)$   // Update parameters using environmental rewards
19:         $\theta\leftarrow\theta+\alpha\nabla\eta_{i}(\theta)$     // Update parameters using guidance rewards
20:     end if
21:end for
@@ Appendix D Proof of Lemma 1
Proof: Consider the environments with sparse rewards, the agent will not receive any non-zero reward unless it reaches the specific goals. Hence, the return of any trajectory generated by the agent is given as:
@@
According to the definition of the trajectory weight function in Eq. (10) and the joint return function (12), a current trajectory can obtain the maximal smooth guidance reward only when it has the smallest distance from the replay memory $\mathcal{P}$ and the agent finds the optimal reward along this trajectory. The smooth guidance reward in Eq. (13) is the expectation of the product of the weight and joint return value for trajectories through $(s,a)$. Therefore, the smooth guidance reward for each state-action pair $(s,a)$ is maximal only when all the trajectories through $(s,a)$ are closest to $\mathcal{P}$. Moreover, the value of $r_{i}(s,a)$ for the state-action pair in $\mathcal{P}$ is larger than that of the other state-action pair.
@@
Appendix E Proof of Theorem 1 and Corollary 1
@@
$J(\pi^{\prime})-J(\pi)\geq\frac{1}{1-\gamma}\left(T_{\pi,f}(\pi^{\prime})-2\epsilon_{f}^{\pi^{\prime}}D_{TV}(d^{\pi^{\prime}}\parallel d^{\pi})\right),$ (27)
$J(\pi^{\prime})-J(\pi)\leq\frac{1}{1-\gamma}\left(T_{\pi,f}(\pi^{\prime})+2\epsilon_{f}^{\pi^{\prime}}D_{TV}(d^{\pi^{\prime}}\parallel d^{\pi})\right),$ (28)
@@
$\|d^{\pi^{\prime}}-d^{\pi}\|_{1}\leq\frac{2\gamma}{1-\gamma}\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{TV}(\pi^{\prime}\parallel\pi)[s]\right],$ (29)
where $D_{TV}(\pi^{\prime}\parallel\pi)[s]=(1/2)\sum_{a}|\pi^{\prime}(a|s)-\pi(a|s)|$ is the total variational divergence at $s$.
@@
Proof: In the proposed methods, we use environmental rewards and smooth guidance rewards to compute the policy gradient and update the policy parameters in a two-step optimization manner. We can then analyze each optimization step’s policy performance improvement bound individually.
First, in the environmental reward optimization step, the environmental reward is used to compute the policy gradient and update the policy parameters. With the bounds from Lemma 3 and bound the divergence $D_{TV}(d^{\pi^{\prime}}\parallel d^{\pi})$ by Lemma 4, when $f$ is the value function $V_{e}$ calculated by the environmental reward, we can easily obtain the following result:
@@
$\tilde{D}_{\pi}(\pi^{\prime})\doteq\frac{\tilde{T}_{\pi}(\pi^{\prime})}{1-\gamma}-\frac{2\gamma\tilde{\epsilon}_{\pi^{\prime}}}{(1-\gamma)^{2}}\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{TV}(\pi^{\prime}\parallel\pi)[s]\right],$ (33)
With the bounds from Lemma 3 and bound the divergence $D_{TV}(d^{\pi^{\prime}}\parallel d^{\pi})$ by Lemma 4, the following result can be easily obtained:
@@
$\eta(\pi^{\prime})-\eta(\pi)\geq(1+\lambda)D_{\pi}^{-}(\pi^{\prime}).$ (37)
@@
Proof: The result is obtained by combining Pinsker’s inequality in Eq (20) with Theorem 1.

### 10 Apr 2024 v2

@@ 7.3 Results in the Ant Maze
To investigate how effective POSG is across various environments, we evaluated our proposed POSG algorithm on the hard-exploration Ant Maze task. It turns out that our approach significantly outperforms other baseline methods in terms of control performance. As shown in Fig. 6a, only our method can reach the goal and learn the optimal policy in the hard-exploration ant maze. Other baseline methods cannot find the sparse reward even if the demonstrations are provided. This result demonstrates that our POSG can better exploit sparse-reward demonstrations by merging the demonstrations’ distribution information and reward signals compared with other RLfD methods. Fig. 6b shows the changing curves of the MMD distance in the Ant Maze task. Our POSG method can efficiently decrease the distance between the training policy and sparse-reward demonstrations. Hence, POSG learns expert-like behaviors with demonstrations quickly. In contrast, the PPO+D agent cannot effectively learn the behaviors of demonstrations. Due to the sparsity of the reward function, the agent’s policy often deteriorates during the policy optimization process, and the agent always falls near the initial point at the end of training.
Fig. 6: (a) Learning curves of success rate on the Ant Maze task; (b) Learning curves of average ground truth reward.
@@ 7.4 Ablation Analyses
Fig. 7: (a) Learning curves of average return; (b) Learning curves of average ground truth reward.
7.4.1 Demonstrations with different amounts
We selected five demonstration datasets with different amounts for comparison on the SparseHopper task. Notice that the POSG experiment is conducted with six trajectories of 500 state-action pairs as demonstrations. The black dashed line in Fig. 7a represents the expert-level performance. In Fig. 7b, The $y$-value corresponding to the dashed line is the value of the average reward per step learned by the PPO algorithm under the ground-truth reward setting. Fig. 7 presents the result of this ablation experiment. Our POSG algorithm produces nearly identical results with different numbers of sparse-reward demonstrations. This result suggests that our POSG algorithm can achieve impressive performance using only a few demonstrations, even a single trajectory. Hence, our method decreases the requirement of demonstration quantity, which facilitates the application of the algorithm in real-world scenarios.
Fig. 8: (a) Learning curves of average return; (b) Learning curves of average ground truth reward.
7.4.2 Demonstrations with different qualities
We obtained the demonstration datasets with different qualities by generating and storing the trajectory data in the different training phases. The corresponding result is shown in Fig. 8. In Fig. 8a, the black dashed line represents the expert-level performance, and the gray dashed line represents the medium-level performance. The dashed line in Fig. 8b has the same meaning as that in Fig. 7b. The average return of POSG increases as the proportion of high-quality data increases. This result indicates that the quality of demonstrations will significantly affect the performance of POSG and that high-quality demonstration data can contribute to policy optimization to some extent.
@@ 8 Conclusion
This article proposes a simple and efficient algorithm called Policy Optimization with Smooth Guidance (POSG) that leverages a small set of state-only demonstrations (where only state information is included in demonstrations) to make approximate and effective long-term credit assignments while efficiently facilitating exploration. The key idea is that the relative impact of state-action pairs can be indirectly estimated using offline state-only demonstrations. More specifically, we obtain the importance of a trajectory by considering the distributional distance between policies and the returns of the associated trajectories. Then, the smooth guidance reward is computed by smoothly averaging the trajectory importance over the trajectory space. We theoretically analyze the performance improvement caused by smooth guidance rewards and derive a new worst-case lower bound on the performance improvement. We conducted various experiments by benchmarking POSG with several state-of-the-art RL algorithms in four environments. Experimental results demonstrate POSG’s superiority over other baseline approaches in terms of performance and convergence speed.
@@ Appendix A POSG Algorithm Training Process in Discrete Spaces
0:  learning rate $\alpha$, offline sparse-reward demonstration dataset $\mathcal{P}$, policy update frequency $K$
1:  Initialize policy weights $\theta$ of each agent.
2:  // Buffer that stores key-value pairs of a state-action tuple $(\boldsymbol{s},\boldsymbol{a})$ and a list of values of trajectory importance that include $(\boldsymbol{s},\boldsymbol{a})$
3:  Initialize $\mathcal{M}\leftarrow\emptyset\ \forall(\boldsymbol{s},\boldsymbol{a})$
4:  for $\text{episode}\in\{0,\dots,T\}$ do
5:     $\tau\leftarrow\emptyset$    // Store state-action pairs for current episode
6:     $R\leftarrow 0$    // Accumulate rewards for current episode
7:     for each step in $\{1,\dots,T\}$ do
8:        Choose $\boldsymbol{a}$ from $\boldsymbol{s}$ using $\pi_{\theta}$
9:        take action $\boldsymbol{a}$ and observe $\boldsymbol{r}_{e}$ and $\boldsymbol{s}^{\prime}$
10:        $\tau\leftarrow\tau\cup\{(\boldsymbol{s},\boldsymbol{a})\}$; $R\leftarrow R+\boldsymbol{r}_{e}$
11:     end for
12:     Compute the trajectory weight $\omega(\tau|\mathcal{M}_{E})$ and the joint return $R_{j}(\tau)$
13:     Compute the trajectory importance $I(\tau)=\omega(\tau|\mathcal{M}_{E})R_{j}(\tau)$
14:     for each $((\boldsymbol{s},\boldsymbol{a}))$ in $\tau_{e}$ do
15:        $\boldsymbol{r}_{i}=\mathbb{E}_{I\sim\mathcal{M}(\boldsymbol{s},\boldsymbol{a})}[I]$  // Compute guidance rewards for each state-action pair
16:     end for
17:     if $\text{episode}\,\%\,K==0$ then
18:        Compute $A_{e}$ using environmental rewards
19:        Compute $A_{i}$ using guidance rewards
20:        $\theta\leftarrow\theta+\alpha\nabla\eta_{e}(\theta)$   // Update parameters using environmental rewards
21:        $\theta\leftarrow\theta+\alpha\nabla\eta_{i}(\theta)$     // Update parameters using guidance rewards
22:     end if
23:  end for
@@ Appendix B POSG Algorithm Training Process in High-Dimensional Continuous Spaces
0:  learning rate $\alpha$, offline sparse-reward demonstration dataset $\mathcal{P}$, policy update frequency $K$
1:  Initialize policy weights $\theta$ of each agent.
2:  // Buffer that stores key-value pairs of a state-action tuple $(\boldsymbol{s},\boldsymbol{a})$ and a list of values of trajectory importance that include $(\boldsymbol{s},\boldsymbol{a})$
3:  Initialize $\mathcal{M}\leftarrow\emptyset\ \forall(\boldsymbol{s},\boldsymbol{a})$
4:  for $\text{episode}\in\{0,\dots,T\}$ do
5:     $\tau\leftarrow\emptyset$    // Store state-action pairs for current episode
6:     $R\leftarrow 0$    // Accumulate rewards for current episode
7:     for each step in $\{1,\dots,T\}$ do
8:        Choose $\boldsymbol{a}$ from $\boldsymbol{s}$ using $\pi_{\theta}$
9:        take action $\boldsymbol{a}$ and observe $\boldsymbol{r}_{e}$ and $\boldsymbol{s}^{\prime}$
10:        $\tau\leftarrow\tau\cup\{(\boldsymbol{s},\boldsymbol{a})\}$; $R\leftarrow R+\boldsymbol{r}_{e}$
11:     end for
12:     Compute the trajectory weight $\omega(\tau|\mathcal{M}_{E})$ and the joint return $R_{j}(\tau)$
13:     Compute the guidance reward $I(\tau)=\omega(\tau|\mathcal{M}_{E})R_{j}(\tau)$
14:     if $\text{episode}\,\%\,K==0$ then
15:        Compute $A_{e}$ using environmental rewards
16:        Compute $A_{i}$ using guidance rewards
17:        $\theta\leftarrow\theta+\alpha\nabla\eta_{e}(\theta)$   // Update parameters using environmental rewards
18:        $\theta\leftarrow\theta+\alpha\nabla\eta_{i}(\theta)$     // Update parameters using guidance rewards
19:     end if
20:  end for
@@ Appendix C Proof of Lemma 1
Proof.
Consider the environments with sparse rewards, the agent will not receive any non-zero reward unless it reaches the specific goals. Hence, the return of any trajectory generated by the agent is given as:
@@
According to the definition of the trajectory weight function in Eq. (10) and the joint return function (12), a current trajectory can obtain the maximal importance weight only when it has the smallest distance from the replay memory $\mathcal{P}$. The smooth guidance reward in Eq. (13) is the expectation of the product of the importance weight and joint return value for trajectories through $(s,a)$. Therefore, the smooth guidance reward for each state-action pair $(s,a)$ is maximal only when all the trajectories through $(s,a)$ are closest to $\mathcal{P}$. Moreover, the value of $r_{i}(s,a)$ for the state-action pair in $\mathcal{P}$ is larger than that of the other state-action pair.
@@
Appendix D Proof of Theorem 1 and Corollary 1
@@
$J(\pi^{\prime})-J(\pi)\geq\frac{1}{1-\gamma}\left(T_{\pi,f}(\pi^{\prime})-2\epsilon_{f}^{\pi^{\prime}}D_{\rm TV}(d^{\pi^{\prime}}\parallel d^{\pi})\right),$ (27)
$J(\pi^{\prime})-J(\pi)\leq\frac{1}{1-\gamma}\left(T_{\pi,f}(\pi^{\prime})+2\epsilon_{f}^{\pi^{\prime}}D_{\rm TV}(d^{\pi^{\prime}}\parallel d^{\pi})\right),$ (28)
@@
$\|d^{\pi^{\prime}}-d^{\pi}\|_{1}\leq\frac{2\gamma}{1-\gamma}\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{\rm TV}(\pi^{\prime}\parallel\pi)[s]\right],$ (29)
where $D_{\rm TV}(\pi^{\prime}\parallel\pi)[s]=(1/2)\sum_{a}|\pi^{\prime}(a|s)-\pi(a|s)|$ is the total variational divergence at $s$.
@@
Proof.
In the proposed methods, we use environmental rewards and smooth guidance rewards to compute the policy gradient and update the policy parameters in a two-step optimization manner. We can then analyze each optimization step’s policy performance improvement bound individually.
First, in the environmental reward optimization step, the environmental reward is used to compute the policy gradient and update the policy parameters. With the bounds from Lemma 3 and bound the divergence $D_{\rm TV}(d^{\pi^{\prime}}\parallel d^{\pi})$ by Lemma 4, when $f$ is the value function $V_{e}$ calculated by the environmental reward, we can easily obtain the following result:
@@
$\tilde{D}_{\pi}(\pi^{\prime})\doteq\frac{\tilde{T}_{\pi}(\pi^{\prime})}{1-\gamma}-\frac{2\gamma\tilde{\epsilon}_{\pi^{\prime}}}{(1-\gamma)^{2}}\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{\rm TV}(\pi^{\prime}\parallel\pi)[s]\right],$ (33)
With the bounds from Lemma 3 and bound the divergence $D_{\rm TV}(d^{\pi^{\prime}}\parallel d^{\pi})$ by Lemma 4, the following result can be easily obtained:
@@
$\eta(\pi^{\prime})-\eta(\pi)\geq(1+\lambda)D_{\pi}(\pi^{\prime}).$ (37)
@@
Proof.
The result is obtained by combining Pinsker’s inequality in Eq (20) with Theorem 1. ∎
Acknowledgment
This work was supported by the National Key R&D Program of China (2022ZD0116401), Fundamental Research Funds for the Central Universities and the National Natural Science Foundation of China(Grant Nos. 62141605)
Conflict of interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper
Authors’ Contributions
Guojian Wang: Conceptualization, Methodology, Software, Formal analysis, Writing–Original Draft. Faguo Wu: Writing–Review and Editing, Supervision. Xiao Zhang: Validation, Supervision, Funding acquisition. Tianyuan Chen: Formal analysis, Software, Visualization. Zhiming Zheng: Validation, Supervision, Funding acquisition.
Availability of data and materials
The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.